# Insights into the Production of 44Ti and Nickel Isotopes in Core-Collapse Supernovae

### v1

@@ 2 Method
Nucleosynthetic yields in this paper are calculated in the same way as in Wang & Burrows (2024a). We use post-processed backwardly-integrated tracer particles to track the dynamical and thermal evolution of the ejected matter. About 320,000 tracers are injected into each simulation. The tracers are placed logarithmically along the r-direction above 500 km and uniformly along the $\theta$- and $\phi$-directions at the end of each simulation. A more detailed description of the tracer method can be found in Wang & Burrows (2024a). The tracer trajectories with neutrino field information are fed into SkyNet (Lippuner & Roberts, 2017), and a 1530-isotope network including elements up to $A=100$ is used to calculate the nucleosynthetic results. The reaction rates are taken from the JINA Reaclib (Cyburt et al., 2010) database, and we include neutrino interactions with protons and neutrons. However, reactions for the $\nu-$process (Woosley et al., 1990) are not included. The NSE (nuclear statistical equilibrium) criterion is set at 0.6 MeV ($\sim$7 GK). The electron fractions ($Y_{e}$) are calculated by Fornax when the temperature is above the NSE threshold, which allows the neutrino spectra to be appropriately non-thermal. A nucleosynthesis calculation starts from the point after which the tracers never reach NSE again. We use the Lodders (2021) values whenever solar abundances are referred to in this paper.
@@
$M_{\rm ZAMS}$ [$M_{\odot}$] $t_{\rm end}$ [s] $E_{\rm exp}$ [B] $M_{\rm NSE}$ ($M_{\rm extra})$[$M_{\odot}$] 44Ti [$10^{-4}M_{\odot}$] 56Ni [$M_{\odot}$] 57Ni [$10^{-3}M_{\odot}$] Stable Ni/Fe
9(a) 1.775 0.111 0.0128 ($4.88\times 10^{-4}$) 0.013 0.0019 0.061 0.502
9(b) 1.950 0.094 0.0115 ($3.35\times 10^{-4}$) 0.035 0.0063 0.065 0.055
9.25 3.532 0.124 0.0198 ($3.17\times 10^{-4}$) 0.074 0.0106 0.116 0.054
9.5 2.375 0.142 0.0262 ($2.24\times 10^{-3}$) 0.121 0.0156 0.180 0.054
11 4.492 0.321 0.0618 ($3.02\times 10^{-2}$) 0.413 0.0418 0.763 0.216
15.01 4.383 0.288 0.0703 ($5.64\times 10^{-3}$) 0.680 0.0548 1.212 0.056
16 4.184 0.390 0.0823 ($1.24\times 10^{-2}$) 0.648 0.0653 1.025 0.045
17 6.164 1.123 0.1889 ($7.63\times 10^{-3}$) 1.443 0.1123 2.706 0.079
18 8.617 0.516 0.1480 ($2.44\times 10^{-2}$) 1.055 0.1181 2.219 0.054
18.5 6.408 1.155 0.1760 ($8.96\times 10^{-3}$) 1.192 0.1421 2.335 0.043
19 4.075 0.466 0.1118 ($7.06\times 10^{-2}$) 1.095 0.0949 1.515 0.052
19.56 3.890 2.451 0.3203 (0) 1.619 0.2536 3.594 0.047
20 6.429 0.881 0.1437 ($1.28\times 10^{-2}$) 1.080 0.1087 2.126 0.034
23 6.228 0.513 0.1223 ($1.92\times 10^{-2}$) 1.180 0.0944 1.627 0.055
24 3.919 0.779 0.1586 ($2.64\times 10^{-2}$) 1.287 0.1348 2.155 0.039
25 6.419 1.101 0.2419 ($3.18\times 10^{-2}$) 2.124 0.1887 3.128 0.037
40 1.760 1.804 0.2151 (0) 1.162 0.1798 4.661 0.106
60 8.026 0.541 0.1616 ($4.82\times 10^{-2}$) 1.645 0.1321 2.785 0.117
Table 1: This table summarizes some basic properties of our models. The core bounce time is set to $t=0$. $t_{\rm end}$ is the time after bounce when the full 3D simulation stops, $E_{\rm exp}$ is the explosion energy in Bethes ($\equiv$1051 ergs), $M_{\rm NSE}$ is the mass of ejecta that has ever reached our NSE threshold (0.6 MeV), and $M_{\rm extra}$ is the mass of the extrapolated neutrino-driven winds. The isotopic yields are calculated at about twenty seconds post-bounce. The solar stable Ni/Fe ratio is 0.058 (Lodders, 2021), and the ratios in our models (including the residual progenitor envelope) range between about 0.6 to 7 times the solar value. The 19.56 and 40 $M_{\odot}$ models form black holes at the end of the simulations (Burrows et al., 2023, 2024); thus, they have $M_{\rm extra}=0$.
@@ 3 Results
The production of 44Ti is different from that of 56Ni. Magkotsios (2011) has studied in detail the production channels of 44Ti, and we have used this work to identify the most important channel for 44Ti production. We find that the dominant contribution to 44Ti production in our simulations comes from the neutrino-driven winds and corresponds to the “(p,$\gamma$)-leakage” regime discussed in Magkotsios (2011). This is a special type of freeze-out. Briefly, in a proton-rich environment (which is very common in neutrino-driven winds), the (p,$\gamma$) reactions transfer the nucleosynthetic flow from symmetric to proton-rich isotopes. As the temperature decreases, the freeze-out process starts and the weak interactions decrease $Y_{e}$ and cause the flow to transfer back towards more stable isotopes. In the case of 44Ti, this means that isotopes such as 45V and 46Cr are produced at early times, and they are converted into 44Ti as the temperature decreases. This explains the non-monotonic growth of the 44Ti abundances in our tracers. The decay of 44V to 44Ti also contributes to 44Ti growth, but plays only a secondary role. The freeze-out process remains active even when the temperature is lower than 1 GK. Therefore, the long production timescale of 44Ti (Sieverding et al., 2023a; Wang & Burrows, 2024a) is due to two main factors: (a) the neutrino-driven winds last for many seconds and keep contributing to the 44Ti yield and (b) the freeze-out process requires a very low temperature to reach the asymptotic state, and this takes a long time to achieve during the expansion. Figure 3 shows the 44Ti yields and 44Ti/56Ni ratios of all our models as a function of time. The circular dot on each curve marks the point when the extrapolation is started. Many models take more than 10 seconds to reach their asymptotic 44Ti yields, and $10^{-4}M_{\odot}$ of 44Ti production can easily be achieved by models more massive than $\sim$16 $M_{\odot}$. More massive models in general produce more 44Ti (see Table 1). The right panel shows that the more massive models also tend to have higher 44Ti/56Ni ratios than the low mass models (e.g., 9a, 9b, 9.25, and 9.5 $M_{\odot}$), so the more massive models not only have higher ejecta masses but also a higher 44Ti production efficiency. This is because more massive explosion models are in general more asymmetric than the low-mass progenitor models due to their generally more massive envelopes.
@@
Multi-dimensional effects play a central role in the production of 44Ti. As Sieverding et al. (2023a) has pointed out, large-scale convection leads to non-monotonic thermal histories. Thus, nucleosynthesis can experience multiple nucleosynthetic phases, which aid the production of 44Ti. We see the same behavior in our 3D simulations. However, this effect can’t solely explain the significantly higher 44Ti yields in our simulations, since the ejecta with significantly non-monotonic thermal histories contribute only a small fraction to the 44Ti yields. The most important 3D effect is the so-called “simultaneous explosion and accretion” phenomenon. This phenomenon is sometimes referred to as “long-lasting” or “fallback” accretion. In the context of this phenomenon, the stellar envelope above the proto-neutron star (PNS) is not immediately swept away by the shock wave after the launch of the explosion. Instead, it forms a few inward-directed funnels that extend almost to the PNS surface. Matter in the outer layers can fall very close to the PNS through these funnels, after which it is heated by neutrinos and ejected as part of the anisotropic neutrino-driven wind. The simultaneous explosion and accretion phenomenon significantly increases the amount of mass interacting with neutrinos, which adds to the $\alpha$-rich freeze-out component. This enhancement of $\alpha$-rich freeze-out lasts for many seconds, until accretion is completely terminated and a spherical wind develops [^1] For examples of spherical winds see Stockinger et al. (2020) and Wang & Burrows (2024b). In addition, interactions with neutrinos push $Y_{e}$ to higher values, which provides the proton-rich environment required by the “(p,$\gamma$)-leakage” channel. Therefore, the simultaneous explosion and accretion phenomenon (redirecting fallback into a component of the late-time winds) aids the production of 44Ti by enhancing the proton-rich neutrino-driven winds. This is the origin of the high 44Ti yields we observed in our long-term 3D CCSN simulations.
@@
Figure 4 plots the 44Ti and 56Ni yields predicted by our 3D theoretical CCSN models, along with the observed Cas A and SN1987A yields. The observed 44Ti and 56Ni masses of Cas A are taken from Grefenstette et al. (2017) and Hwang & Laming (2012), while the SN1987A values are taken from Boggs et al. (2015) and McCray & Fransson (2016). Our 3D models can easily produce high amounts of 44Ti and reach the high 44Ti/56Ni ratio observed in Cas A. In addition, the lower 44Ti/56Ni ratios of the black-hole formers clearly show the important role of neutrino-driven winds in the production of 44Ti, since the winds in these models are turned off after the black hole formation. Therefore, the “44Ti-problem” (The et al., 2006) is no longer an issue in self-consistent long-term 3D initially non-rotating CCSN models. SN1987A is still about two sigmas above the 44Ti/56Ni ratios predicted by our models. We now turn to a discussion of this issue and its corresponding 57Ni/56Ni ratio.
@@ 3.2 Effects of the Extrapolation Method
Since a significant amount of 44Ti and 57Ni is produced during the extrapolated phase, will the above results be changed by the choice of extrapolation method? The contribution from the extrapolated neutrino-driven wind accounts for no more than 20% of the yield of these radioactive isotopes, so the wind extrapolation methods should not influence the results to any significant degree. However, it is not impossible that how we chose to extrapolate the tracer thermal histories might change our conclusions, since this influences most of the ejecta that never reaches high temperatures. Therefore, it would be helpful to have a robustness test of our choice of tracer extrapolation method.
@@
where $t_{\rm end}$ is the post-bounce duration of the 3D simulations and $\tau$ is an expansion timescale fitted using the last 10 ms of the trajectories. The left panel of Figure 7 compares the power-law and exponential methods using an example tracer. The exponential method leads to a much faster decrease of temperature and density than the power-law method. Therefore, nucleosynthesis will terminate significantly earlier if the exponential method is used. We apply the exponential method to the 25 $M_{\odot}$ model, whose power-law version manifests one of the greatest 44Ti and 57Ni yield growths during the extrapolated phase. The results are shown in the right panel of Figure 7. We find that the choice of extrapolation methods can change the final yield by about 20%. This is not a negligible influence, but a $\sim$20uncertainty is tolerable in the context of current CCSN nucleosynthesis studies and does not change our general conclusions. The impact of the choice of extrapolation method is weaker for other models (e.g. the 9.x models, 17, 18, and 60 models), because these 3D simulations themselves have already captured much of the late-time yield growth.
@@
Figure 7: Comparison of the two thermal history extrapolation methods and their influence on the nucleosynthetic results. Blue lines show the evolution before extrapolation, green lines show the power-law extrapolation, and red lines show the exponential extrapolation. Left: Temperature and density evolution of an example tracer selected from the 25 $M_{\odot}$ model. The temperature of the tracer is about 3 GK at the beginning of the extrapolation. Right: Total 44Ti and 57Ni produced in the 25 $M_{\odot}$ model as a function of time. It’s clear that the power-law extrapolation allows the tracer to stay at relatively higher temperature for a longer period of time, and thus enhances the yields. Different extrapolation methods lead to about 20% deviation in the final isotopic yields of the 25 $M_{\odot}$ model, which is the model most sensitive to extrapolation choices.
3.3 Stable Isotopes of Nickel and Stable Nickel to Iron Ratios

### v2

@@ 2 Method
Nucleosynthetic yields in this paper are calculated in the same way as in Wang & Burrows (2024a). We use post-processed backwardly-integrated tracer particles to track the dynamical and thermal evolution of the ejected matter. About 320,000 tracers are injected into each simulation. The tracers are placed logarithmically along the r-direction above 500 km and uniformly along the $\theta$- and $\phi$-directions at the end of each simulation. A more detailed description of the tracer method can be found in Wang & Burrows (2024a). The tracer trajectories with neutrino field information are fed into SkyNet (Lippuner & Roberts, 2017), and a 1530-isotope network including elements up to $A=100$ is used to calculate the nucleosynthetic results. The reaction rates are taken from the JINA Reaclib (Cyburt et al., 2010) database, and we include neutrino interactions with protons and neutrons. However, reactions for the $\nu-$process (Woosley et al., 1990) are not included. The NSE (nuclear statistical equilibrium) criterion is set at 0.6 MeV ($\sim$7 GK). The electron fractions ($Y_{e}$) are calculated by Fornax when the temperature is above the NSE threshold, which allows the neutrino spectra to be appropriately non-thermal. A nucleosynthesis calculation starts from the point after which the tracers never reach NSE again. We use the Lodders (2021) values whenever solar abundances are referred to in this paper. If not otherwise stated, all ratios mentioned in the paper are mass fraction ratios.
@@
$M_{\rm ZAMS}$ $t_{\rm end}$ $E_{\rm exp}$ $M^{\rm pow}_{{}^{44}{\rm Ti}}$($M^{\rm exp}_{{}^{44}{\rm Ti}}$) $M^{\rm pow}_{{}^{56}{\rm Ni}}$($M^{\rm exp}_{{}^{56}{\rm Ni}}$) $M^{\rm pow}_{{}^{57}{\rm Ni}}$($M^{\rm exp}_{{}^{57}{\rm Ni}}$) Stable Ni/Fe $M_{\rm extra}$
$[M_{\odot}]$ [s] [B] [$10^{-4}M_{\odot}$] [$M_{\odot}$] [$10^{-3}M_{\odot}$] [$M_{\odot}$]
9(a) 1.775 0.111 0.013(0.013) 0.0019(0.0019) 0.061(0.061) 0.502 $4.88\times 10^{-4}$
9(b) 1.950 0.094 0.035(0.034) 0.0063(0.0062) 0.065(0.064) 0.055 $3.35\times 10^{-4}$
9.25 3.532 0.124 0.074(0.074) 0.0106(0.0106) 0.116(0.116) 0.054 $3.17\times 10^{-4}$
9.5 2.375 0.142 0.121(0.107) 0.0160(0.0160) 0.180(0.165) 0.054 $2.24\times 10^{-3}$
11 4.492 0.321 0.413(0.408) 0.0418(0.0418) 0.763(0.759) 0.216 $3.02\times 10^{-2}$
15.01 4.383 0.288 0.680(0.566) 0.0548(0.0548) 1.212(0.931) 0.056 $5.64\times 10^{-3}$
16 4.184 0.390 0.648(0.565) 0.0653(0.0655) 1.025(0.850) 0.045 $1.24\times 10^{-2}$
17 6.164 1.123 1.443(1.392) 0.1123(0.1124) 2.706(2.611) 0.079 $7.63\times 10^{-3}$
18 8.617 0.516 1.055(1.032) 0.1181(0.1182) 2.219(2.157) 0.054 $2.44\times 10^{-2}$
18.5 6.408 1.155 1.192(1.064) 0.1421(0.1426) 2.335(1.993) 0.043 $8.96\times 10^{-3}$
19 4.075 0.466 1.095(0.945) 0.0949(0.0955) 1.515(1.188) 0.052 $7.06\times 10^{-2}$
19.56 3.890 2.451 1.619(1.284) 0.2536(0.2550) 3.594(2.706) 0.047 0
20 6.429 0.881 1.080(0.972) 0.1087(0.1091) 2.126(1.832) 0.034 $1.28\times 10^{-2}$
23 6.228 0.513 1.180(1.044) 0.0944(0.0948) 1.627(1.346) 0.055 $1.92\times 10^{-2}$
24 3.919 0.779 1.287(1.110) 0.1348(0.1359) 2.155(1.525) 0.039 $2.64\times 10^{-2}$
25 6.419 1.101 2.124(1.838) 0.1887(0.1899) 3.128(2.491) 0.037 $3.18\times 10^{-2}$
40 1.760 1.804 1.162(0.935) 0.1798(0.1817) 4.661(3.661) 0.106 0
60 8.026 0.541 1.645(1.572) 0.1321(0.1323) 2.785(2.598) 0.117 $4.82\times 10^{-2}$
Table 1: This table summarizes some basic properties of our models. The core bounce time is set to $t=0$. $t_{\rm end}$ is the time after bounce when the full 3D simulation stops, $E_{\rm exp}$ is the explosion energy in Bethes ($\equiv$1051 ergs). The isotopic yields are calculated at about twenty seconds post-bounce. $M^{\rm pow}$ is the mass calculated using the power-law tracer extrapolation method, while $M^{\rm exp}$ is calculated using the exponential tracer extrapolation method. Discussions and plots in this paper are based on the power-law results, while the exponential results are listed to indicate that there can be up to a $\sim$20uncertainty in certain isotope yields. A more detailed discussion on the uncertainties can be found in Section 3.3. The solar stable Ni/Fe ratio is 0.058 (Lodders, 2021), and the ratios in our models (including the residual progenitor envelope) range between about 0.6 to 7 times the solar value. $M_{\rm extra}$ is the mass of the neutrino-driven winds in the extrapolated phases. The 19.56 and 40 $M_{\odot}$ models form black holes at the end of the simulations (Burrows et al., 2023, 2024); thus, they have $M_{\rm extra}=0$.
@@ 3 Results
The production of 44Ti is different from that of 56Ni. Magkotsios (2011) has studied in detail the production channels of 44Ti, and we have used this work to identify the most important channel for 44Ti production. We find that the dominant contribution to 44Ti production in most of our simulations [^1] Model 9(a) is a special case because it has a strong neutron-rich phase. This phase suppresses the (p,$\gamma$)-leakage channel at early times and leads to a smaller amount of 44Ti. But because 56Ni is more suppressed, the 44Ti/56Ni ratio in 9(a) is higher than that in the more proton-rich 9(b). comes from the neutrino-driven winds and corresponds to the “(p,$\gamma$)-leakage” regime discussed in Magkotsios (2011). This is a special type of freeze-out. Briefly, in a proton-rich environment (which is very common in neutrino-driven winds), the (p,$\gamma$) reactions transfer the nucleosynthetic flow from symmetric to proton-rich isotopes. As the temperature decreases, the freeze-out process starts and the weak interactions decrease $Y_{e}$ and cause the flow to transfer back towards more stable isotopes. In the case of 44Ti, this means that isotopes such as 45V and 46Cr are produced at early times, and they are converted into 44Ti as the temperature decreases. This explains the non-monotonic growth of the 44Ti abundances in our tracers. The decay of 44V to 44Ti also contributes to 44Ti growth, but plays only a secondary role. The freeze-out process remains active even when the temperature is lower than 1 GK because of the ongoing weak interactions. Therefore, the long production timescale of 44Ti (Sieverding et al., 2023a; Wang & Burrows, 2024a) is due to two main factors: (a) the neutrino-driven winds last for many seconds and keep contributing to the 44Ti yield and (b) the freeze-out process requires a very low temperature to reach the asymptotic state, and this takes a long time to achieve during the expansion. Figure 3 shows the 44Ti yields and 44Ti/56Ni ratios of all our models as a function of time. The circular dot on each curve marks the point when the extrapolation is started. Many models take more than 10 seconds to reach their asymptotic 44Ti yields, and $10^{-4}M_{\odot}$ of 44Ti production can easily be achieved by models more massive than $\sim$16 $M_{\odot}$. More massive models in general produce more 44Ti (see Table 1). The right panel shows that the more massive models also tend to have higher 44Ti/56Ni ratios than the low mass models (e.g., 9a, 9b, 9.25, and 9.5 $M_{\odot}$), so the more massive models not only have higher ejecta masses but also a higher 44Ti production efficiency. This is because more massive explosion models are in general more asymmetric than the low-mass progenitor models due to their generally more massive envelopes.
@@
Explosion asymmetry and general multi-dimensional effects play central roles in the production of 44Ti. As Sieverding et al. (2023a) has pointed out, large-scale convection leads to non-monotonic thermal histories. Thus, nucleosynthesis can experience multiple nucleosynthetic phases, which aid the production of 44Ti. We see the same behavior in our 3D simulations. However, this effect can’t solely explain the significantly higher 44Ti yields in our simulations, since the ejecta with significantly non-monotonic thermal histories contribute only a small fraction to the 44Ti yields. The most important 3D effect is the so-called “simultaneous explosion and accretion” phenomenon. This phenomenon is sometimes referred to as “long-lasting” or “fallback” accretion. In the context of this phenomenon, the stellar envelope above the proto-neutron star (PNS) is not immediately swept away by the shock wave after the launch of the explosion. Instead, it forms a few inward-directed funnels that extend almost to the PNS surface. Matter in the outer layers can fall very close to the PNS through these funnels, after which it is heated by neutrinos and ejected as part of the anisotropic neutrino-driven wind. The simultaneous explosion and accretion phenomenon significantly increases the amount of mass interacting with neutrinos, which adds to the $\alpha$-rich freeze-out component. This enhancement of $\alpha$-rich freeze-out lasts for many seconds, until accretion is completely terminated and a spherical wind develops [^2] For examples of spherical winds see Stockinger et al. (2020) and Wang & Burrows (2024b). In addition, interactions with neutrinos push $Y_{e}$ to higher values, which provides the proton-rich environment required by the “(p,$\gamma$)-leakage” channel. Therefore, the simultaneous explosion and accretion phenomenon (redirecting fallback into a component of the late-time winds) aids the production of 44Ti by enhancing the proton-rich neutrino-driven winds. This is the origin of the high 44Ti yields we observed in our long-term 3D CCSN simulations. Hence, more asymmetric explosions naturally lead to more vigorous simultaneous explosion and accretion behavior and, thus, to higher 44Ti production.
@@
Figure 4 plots the 44Ti and 56Ni yields predicted by our 3D theoretical CCSN models, along with the observed Cas A and SN1987A yields. The observed 44Ti and 56Ni masses of Cas A are taken from Grefenstette et al. (2017) and Hwang & Laming (2012), while the SN1987A values are taken from Boggs et al. (2015) and McCray & Fransson (2016). Our 3D models can easily produce high amounts of 44Ti and reach the high 44Ti/56Ni ratio observed in Cas A. The possible presence of unshocked Fe-rich material in Cas A may move the data point rightward, which would be even more consistent with our predictions. In addition, the lower 44Ti/56Ni ratios of the black-hole formers clearly show the important role of neutrino-driven winds in the production of 44Ti, since the winds in these models are turned off after the black hole formation. Therefore, the “44Ti-problem” (The et al., 2006) is no longer an issue in self-consistent long-term 3D initially non-rotating CCSN models. SN1987A is still about two sigmas above the 44Ti/56Ni ratios predicted by our models. We now turn to a discussion of this issue and its corresponding 57Ni/56Ni ratio.
@@
3.2 Stable Isotopes of Nickel and Stable Nickel to Iron Ratios
@@ 3.3 Effects of the Extrapolation Method
Since a significant amount of 44Ti and 57Ni is produced during the extrapolated phase, will the above results be changed by the choice of extrapolation method? The contribution from the extrapolated neutrino-driven winds accounts for no more than 20% of the yield of these radioactive isotopes [^6] Neutrino-driven winds start shortly after the explosion (Wang & Burrows, 2023), and the first few seconds of wind evolution before the extrapolated phase have been captured by our 3D simulations. Hence, since the wind mass outflow rates decay approximately exponentially with time (Figure 1), the bulk of the wind mass ejection occurs before extrapolation. What is left is the long-tail contribution from neutrino-driven winds during the extrapolated phase, whose estimated masses are listed as $M_{\rm extra}$ in Table 1. The contribution to each individual isotope is then calculated by multiplying $M_{\rm extra}$ with the corresponding isotopic mass fractions at late time: 56Nisimilar-$\sim$0.45, 44Tisimilar-$\sim$0.0004, and 57Nisimilar-$\sim$0.004. Given the small $M_{\rm extra}$, the extrapolated winds can not have a strong impact on the final nucleosynthetic results., so the wind extrapolation does not influence the results to any significant degree. However, it is not impossible that how we chose to extrapolate the tracer thermal histories might change our conclusions, since this influences most of the ejecta that never reaches high temperatures. Therefore, it would be helpful to have a robustness test of our choice of tracer extrapolation method.
@@
where $t_{\rm end}$ is the post-bounce duration of the 3D simulations and $\tau$ is an expansion timescale fitted using the last 10 ms of the trajectories. The exponential method leads to a faster decrease of temperature and density than the power-law method. Therefore, nucleosynthesis will terminate significantly earlier if the exponential method is used.
The results for the exponential method applied to all models are listed in Table 1. The uncertainty of the 56Ni yield due to extrapolation method is $<$1%, while it’s $\sim$10for 44Ti and $\sim$20for 57Ni. The different sensitivities to the extrapolation method for these three isotopes are directly related to their different production timescales. Isotopes with longer production timescales will be more dependent upon the choice of extrapolation method. A $\sim$20uncertainty is not negligible, but it is tolerable in the context of current CCSN nucleosynthesis studies and does not change our general conclusions. The impact of the choice of extrapolation method is weaker in some models(e.g. the 9.x models, 17, 18, and 60 models), as they have already captured much of the late-time yield growth.